\begin{thebibliography}{10}

\bibitem{AyoubThesis}
Joseph Ayoub.
\newblock Les six op\'{e}rations de {G}rothendieck et le formalisme des cycles
  \'{e}vanescents dans le monde motivique. {(I, II)}.
\newblock {\em Ast\'{e}risque}, (314, 315), 2007.

\bibitem{Brokemper}
Dennis Brokemper.
\newblock On the {C}how ring of the stack of truncated {B}arsotti-{T}ate
  groups.
\newblock {\em Pacific J. Math.}, 296(2):271--303, 2018.
\newblock \href {https://doi.org/10.2140/pjm.2018.296.271}
  {\path{doi:10.2140/pjm.2018.296.271}}.

\bibitem{CD}
Denis-Charles Cisinski and Fr\'{e}d\'{e}ric D\'{e}glise.
\newblock {\em Triangulated categories of mixed motives}.
\newblock Springer Monographs in Mathematics. Springer, Cham, [2019] \copyright
  2019.
\newblock \href {https://doi.org/10.1007/978-3-030-33242-6}
  {\path{doi:10.1007/978-3-030-33242-6}}.

\bibitem{HV}
Urs Hartl and Eva Viehmann.
\newblock The {N}ewton stratification on deformations of local {$G$}-shtukas.
\newblock {\em J. Reine Angew. Math.}, 656:87--129, 2011.
\newblock \href {https://doi.org/10.1515/CRELLE.2011.044}
  {\path{doi:10.1515/CRELLE.2011.044}}.

\bibitem{Hoyois}
Marc Hoyois.
\newblock A quadratic refinement of the {G}rothendieck-{L}efschetz-{V}erdier
  trace formula.
\newblock {\em Algebr. Geom. Topol.}, 14(6):3603--3658, 2014.
\newblock \href {https://doi.org/10.2140/agt.2014.14.3603}
  {\path{doi:10.2140/agt.2014.14.3603}}.

\bibitem{LauDisp}
Eike Lau.
\newblock Smoothness of the truncated display functor.
\newblock {\em J. Amer. Math. Soc.}, 26(1):129--165, 2013.
\newblock \href {https://doi.org/10.1090/S0894-0347-2012-00744-9}
  {\path{doi:10.1090/S0894-0347-2012-00744-9}}.

\bibitem{Levine-t}
Marc Levine.
\newblock Tate motives and the vanishing conjectures for algebraic
  {$K$}-theory.
\newblock In {\em Algebraic {$K$}-theory and algebraic topology ({L}ake
  {L}ouise, {AB}, 1991)}, volume 407 of {\em NATO Adv. Sci. Inst. Ser. C: Math.
  Phys. Sci.}, pages 167--188. Kluwer Acad. Publ., Dordrecht, 1993.
\newblock URL: \url{https://doi.org/10.1007/978-94-017-0695-7_7}, \href
  {https://doi.org/10.1007/978-94-017-0695-7\_7}
  {\path{doi:10.1007/978-94-017-0695-7\_7}}.

\bibitem{PWZ}
Richard Pink, Torsten Wedhorn, and Paul Ziegler.
\newblock {$F$}-zips with additional structure.
\newblock {\em Pacific J. Math.}, 274(1):183--236, 2015.
\newblock \href {https://doi.org/10.2140/pjm.2015.274.183}
  {\path{doi:10.2140/pjm.2015.274.183}}.

\bibitem{RS1}
Timo Richarz and Jakob Scholbach.
\newblock The intersection motive of the moduli stack of shtukas.
\newblock {\em Forum Math. Sigma}, 8:Paper No. e8, 99, 2020.
\newblock \href {https://doi.org/10.1017/fms.2019.32}
  {\path{doi:10.1017/fms.2019.32}}.

\bibitem{Robalo}
Marco Robalo.
\newblock {$K$}-theory and the bridge from motives to noncommutative motives.
\newblock {\em Adv. Math.}, 269:399--550, 2015.
\newblock \href {https://doi.org/10.1016/j.aim.2014.10.011}
  {\path{doi:10.1016/j.aim.2014.10.011}}.

\bibitem{Scholze6ff}
Peter Scholze.
\newblock Six-functor formalisms.
\newblock \url{https://people.mpim-bonn.mpg.de/scholze/SixFunctors.pdf}, 2022.

\bibitem{stacks-project}
The {Stacks Project Authors}.
\newblock \textit{Stacks Project}.
\newblock \url{https://stacks.math.columbia.edu}, 2025.

\bibitem{VW}
Eva Viehmann, Torsten Wedhorn, and Appendix by~Christopher~Lang.
\newblock Moduli of truncated shtukas and displays, 2025.
\newblock URL: \url{https://arxiv.org/abs/2506.01740}, \href
  {http://arxiv.org/abs/2506.01740} {\path{arXiv:2506.01740}}.

\bibitem{Voev}
V.~Voevodsky.
\newblock Homology of schemes.
\newblock {\em Selecta Math. (N.S.)}, 2(1):111--153, 1996.
\newblock \href {https://doi.org/10.1007/BF01587941}
  {\path{doi:10.1007/BF01587941}}.

\bibitem{WZ}
Torsten Wedhorn and Paul Ziegler.
\newblock Tautological rings of {S}himura varieties and cycle classes of
  {E}kedahl-{O}ort strata.
\newblock {\em Algebra Number Theory}, 17(4):923--980, 2023.
\newblock \href {https://doi.org/10.2140/ant.2023.17.923}
  {\path{doi:10.2140/ant.2023.17.923}}.

\bibitem{MotGZip}
Can Yaylali.
\newblock Motivic homotopy theory of the classifying stack of finite groups of
  lie type, 2023.
\newblock \href {http://arxiv.org/abs/2306.09808} {\path{arXiv:2306.09808}}.

\bibitem{MotPro}
Can Yaylali.
\newblock Rational motives on pro-algebraic stacks, 2024.
\newblock \href {http://arxiv.org/abs/2403.03596} {\path{arXiv:2403.03596}}.

\bibitem{Zhu}
Xinwen Zhu.
\newblock Tame categorical local langlands correspondence, 2025.
\newblock URL: \url{https://arxiv.org/abs/2504.07482}, \href
  {http://arxiv.org/abs/2504.07482} {\path{arXiv:2504.07482}}.

\end{thebibliography}

\documentclass[article,11pt,leqno]{article}
\usepackage{articlemacro}

\newcommand{\DM}{\mathrm{DM}}

\newcommand{\PreStk}{\mathrm{PreStk}}

\newcommand{\Sht}{\mathrm{Sht}}
\newcommand{\GSht}{\mathrm{Sht}_{G}}
\newcommand{\nSht}[1]{#1\textup{-}\Sht}

\newcommand{\Disp}{\mathrm{Disp}}
\newcommand{\quot}[1]{\left[#1\right]}

\newcommand{\Zip}[1]{#1\textup{-Zip}}

\date{}
\title{Truncations and the Motive of the Stack of Local $G$-Shtukas}
\author{Can Yaylali}

\begin{document}
\maketitle

\begin{abstract}
	We compute the rational motive of the stack of local $G$-shtukas, for a split reductive group $G$, representing compactly supported cohomology in terms of the motive of the stack of $G$-zips. This result makes explicit use of the truncated version of local $G$-shtukas established by Viehmann-Wedhorn and the theoretical background on the $*$- and $!$-adjunction for pro-systems of algebraic objects established by the author.
\end{abstract}

\thispagestyle{empty}
\tableofcontents

\section*{Introduction}
Shtukas occupy a central position in the Langlands program for function fields: their moduli spaces provide geometric realizations of automorphic forms and were used by L. Lafforgue to establish the global Langlands correspondence for $\GL_n$.  Their \emph{local} analogues were introduced by Hartl–Viehmann \cite{HV} play an important role in the proof of the tame categorical local Langlands correspondence  by Xinwen Zhu \cite{Zhu}.  In a parallel but complementary direction, Richarz–Scholbach have established a motivic form of the geometric Satake equivalence with $\QQ$-coefficients, providing a motivic incarnation of one of the key structures in the Langlands program \cite{RS1}. 

\medskip

In this article we study the \emph{rational motive} of the stack of local $G$-shtukas, where $G$ is a split reductive group. Our goal is to describe its motivic cohomology with compact support in purely motivic terms and, in particular, to relate it to the finite-type stacks of $G$-zips that appear in its stratification. The resulting comparison shows that
the cohomological complexity of the stack of local $G$-shtukas is completely captured by its truncated models.

\medskip

Local $G$-shtukas, as introduced by Hartl–Viehmann, are defined over $\FF_p$ and may be viewed as function-field analogues of $p$-divisible groups.  Their moduli stacks are quotients of ind-schemes by pro-algebraic groups and highly infinite-dimensional, but recent work of Viehmann–Wedhorn has shown that they can be  pproximated by truncated versions whose strata are described by stacks of $G$-zips \cite{VW}.  Building on this insight, we develop a motivic framework in which these approximations can be compared directly on the level of rational motives. 

\medskip 

The key input is a localization sequence for motives on quotients of ind-schemes by pro-algebraic groups with split pro-unipotent kernel. Using recent advances in the theory of motives on ind- and pro-Artin stacks, we establish the existence of a $\ast$-adjunction and compact-object-preserving pushforward for such quotients, together with a localization  formula.  This technical step allows us to compute motivic cohomology with compact support in families of local shtukas.

\begin{intro-theorem}[\protect{\ref{thm.main}}]
\label{intro-thm}
	Let $G$ be a split reductive group and $T$ a split maximal torus contained in $G$. Let $\mu\in X_{*}(T)_{+}$ be a dominant cocharacter and let $\GSht^{\leq\mu}$ denote the closure\footnote{See Section \ref{sec-GSht} for the notation.} of the stack of local $G$-shtukas of type $\mu$. Then 
	\[
		M^{c}(\GSht^{\leq\mu}) \simeq \bigoplus_{\mu'\leq \mu} M^{c}(\Zip{G}^{\mu'})
	\]
	in $\DM(\FF_{p})$.
	In particular, it is a Tate motive.
\end{intro-theorem}

 This expresses the motivic cohomology of the (infinite-dimensional) stack of local $G$-shtukas in terms of the finite-type stacks of $G$-zips arising through its stratification.  In other words, all cohomological information of $\Sht_G$ can already be recovered from its truncated models. Consequently, all information about the motivic cohomology of  $\Sht_G$ is already encoded in the finite-dimensional stacks of $G$-zips, whose motives and Chow rings are  explicitly known \cite{Brokemper,WZ,MotGZip}. 
 
\medskip

Beyond this structural statement, our methods provide general tools for working with motives on ind- and pro-Artin stacks.  In particular, we prove: 

\begin{intro-proposition}[\protect{cf. Section \ref{sec-complement}}]
	Let $f\colon X\rightarrow Y$ be a morphism of ind-schemes (not necessarily of ind-(finite type)) and let $G$ be a pro-algebraic group with split pro-unipotent kernel acting on $X$ and $Y$ such that $f$ is $G$-equivariant. Then there exists a functor $f_{*}\colon \DM(X/G)\to\DM(Y/G)$ preserving compact objects, with the following properties:
	\begin{enumerate}
		\item The functor $f_{*}$ admits a left adjoint $f^{*}$.
		\item If $f$ is ind-proper, then $f_{*}\simeq f_{!}$.
		\item If $f$ is an open immersion, then $f^{!}\simeq f^{*}$.
	\end{enumerate}
	Moreover, one obtains a localization sequence. In case one has also a lower $!$-functor, we also have $!$/$*$-base change and a projection formula. 
\end{intro-proposition}

The techniques developed here provide a general framework for working with motives on ind- and pro-Artin stacks and should be applicable in related contexts, such as the study of Rapoport–Zink spaces or motivic versions of the Satake equivalence.

\subsection*{Acknowledgement}
I want to thank Chirantan Chowdhury and Torsten Wedhorn for various discussions around this topic.

This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) TRR 326 \textit{Geometry and Arithmetic of Uniformized Structures}, project number 444845124.

\section{The stack of local $G$-shtukas}
\label{sec-GSht}

Let us quickly recall the construction of local $G$-shtukas and its truncated counterpart following \cite{VW}. Let us fix a a prime $p>0$ and let $k=\FF_{q}$ be a finite field of characteristic $p>0$. Further, let $G$ be a split reductive group scheme over $k[\![z]\!]$. Note that $G$ admits a model over $k$, which we also denote by $G$. Moreover, we fix a split maximal torus $T$ of $G$ and denote the set of dominant cocharacters of $T$ by $X_{*}(T)_{+}$. 

As always one can define the loop group of $G$
$$
	LG\colon R\mapsto G(R(\!(z)\!)),
$$
as well as the positive loop group of $G$
$$
	L^{+}G\colon R\mapsto G(R[\![z]\!])
$$
as fpqc-sheaves on the category of $k$-schemes.
These are representable by ind-schemes and $L^{+}G$ is a subsheaf of $LG$. Note that $LG = \colim (LG)_{i}$ is an ind-affine ind-scheme but not of ind-(finite type).

\begin{defi}[\protect{\cite{HV}}]
	A local $G$-shtuka over an $k$-scheme $S$ is a tuple $(\Ecal,\phi)$, where $\Ecal$ is an $L^{+}G$-torsor on $S$ and $\phi\colon \sigma^{*}\Ecal[1/z]\to \Ecal[1/z]$ is an isomorphism of $LG$-torsors\footnote{Here we $\Ecal[1/z]$ denotes the image of $\Ecal$ under the map $\check{H}^{1}(S,L^{+}G)\to \check{H}^{1}(S,LG)$, cf. \cite{HV}.}.
	
	We denote by $\GSht(S)$ the groupoid of local $G$-shtukas over $S$ and by $\GSht$ the associated fpqc-stack.  
\end{defi}

One can realize $\GSht$ as a ind-pro-Artin stack in the following sense. Let $L^{+}G$ act on $LG$ via Frobenius conjugation, i.e. $\Ad_{\sigma}\colon (a,x) \mapsto ax\sigma(a)^{-1}$. Then we see that $\GSht\simeq LG/\Ad_{\sigma} L^{+}G$ as fpqc quotient stacks. In particular, $\Sht_{G}$ admits a covering by $LG$, an ind-scheme. 

For any dominant cocharacter $\mu\in X_{*}(T)_{+}$, the Bruhat stratification on the affine Grassmannian induces a stratification on $LG$. We denote the associated Schubert cells by $LG_{\mu}$ and the associated Schubert varieties by $LG_{\leq \mu}$. 
Note that $LG_{\leq \mu}\subseteq LG$ is a closed embedding and thus $LG_{\leq \mu}$ is a representable by an affine scheme. Moreover, $LG_{\mu}\subseteq LG_{\leq \mu}$ is affine open embedding. 
This defines a stratification on $\coprod_{\mu\in X_{*}(T)_{+}}LG_{\mu}\to LG$.
We denote the associated quotients 
\[
	\GSht^{\leq \mu}\coloneqq \quot{LG_{\leq \mu}/\Ad_{\sigma}L^{+}G}_{\mathrm{fpqc}},\quad \GSht^{\mu}\coloneqq \quot{LG_{\mu}/\Ad_{\sigma}L^{+}G}_{\mathrm{fpqc}}.
\]

Now let us recall some theory on the moduli of $N$-truncated local $G$-shtukas, for $1\leq N\leq \infty$, following \cite{VW}. We will give a definition of this stack by quotients of truncated loop groups. Let us define those first.

For $1\leq N\leq \infty$, we set
$$
	L^{(N)}G : R\mapsto G(R[\![z]\!]/(z^{N})).
$$
Moreover, for any cocharacter $\mu\in X_{*}(T)_{+}$, we obtain groups $E_{N}(G,\mu)$ acting on $L^{(N)}G$ and transition maps $E_{N+1}(G,\mu)\to E_{N}(G,\mu)$ (cf. \cite{VW}). We will not make the group $E_{N}(G,\mu)$ explicit except for $N=1$ later on, since otherwise we will not need the explicit description of $E_{N}(G,\mu)$.

\begin{defi}[\protect{\cite{VW}}]
	We define the moduli stack of $N$-truncated local $G$-shtukas via 
	$$
		\nSht{N}_{G}^{\mu}\coloneqq \quot{L^{(N)}G/E_{N}(G,\mu)}_{\mathrm{fpqc}}.
	$$
\end{defi}
The composition of the canonical map $L^{(N+1)}G\to L^{(N)}G$ as well as the truncation map $E_{N+1}(G,\mu)\to E_{N}(G,\mu)$ yields a truncation map

\[
	\tau_{N}\colon \nSht{(N+1)}_{G}^{\mu}\to \quot{L^{(N)}G/E_{N+1}(G,\mu)} \to \nSht{N}_{G}^{\mu},
\]
where $E_{N+1}(G,\mu)$ acts on $L^{(N)}G$ via restriction along the truncation map.

\begin{rem}
\label{rem-Sht-coprod}
	Similarly to $\GSht$ one can define $\nSht{N}_{G}$ for any $0\leq N\leq \infty$ equipped with a canonical map $\nSht{\infty}_{G}\to\GSht$, cf. \cite{VW}. However, if $G$ is split, we have an equivalence 
	\[
		\nSht{N}_{G}\simeq \coprod_{\mu\in X_{*}(T)_{+}}\nSht{N}_{G}^{\mu}.
	\]
	In particular, we will not recall the specific definition of $\nSht{N}_{G}$ and define it as the coproduct above.
\end{rem}

In this article, we will be interested in pro-algebraic group schemes arising through successive extensions of unipotent group schemes. Let us define this notion.

\begin{defi}
	Let $G$ be a pro-algebraic group indexed by $\NN$. We say that $G$ has \textit{split pro-unipotent kernel} if $\Ker(G\to G_{1})$ is split pro-unipotent, cf. \cite[App. A]{RS1}.
\end{defi}

\begin{rem}
	Let us  note that in the definition of $\nSht{N}_{G}^{\mu}$ as well as $\GSht$ and $\GSht^{\mu}$, we can equivalently consider the \'etale sheafification since $L^{+}G$ and $E_{\infty}(G,\mu)$ have split pro-unipotent kernel \cite[Prop. A.4.9]{RS1}. In particular, we will only consider \'etale stack quotients from here on.
\end{rem}

\begin{rem}
\label{rem-infty}
	Let us remark that the stacks $\nSht{\infty}_{G}^{\mu}\cong \GSht^{\mu}$ are equivalent through a canonical map $\nSht{\infty}_{G}\to \GSht$ \cite[Thm. 6.21 (2)]{VW}. This is the main technical point, which allows us to compute $M^{c}(\Sht_{G}^{\leq \mu})$ (see Definition \ref{def-motive}) in terms of motives of $\Zip{G}^{\mu'}$ for $\mu'\leq \mu$.
\end{rem}

The stack $\nSht{1}_{G}^{\mu}$ can be described more explicitly in terms of $G$-zips. To define the stack of $G$-zips let us note that the cocharacter $\mu$ yields a parabolic $P_{+}\subseteq G$ with Levi-decomposition $P_{+} = U_{+} \rtimes L$. Let $P_{-}$ denote the opposite parabolic with opposite Levi-decomposition $P_{-} = U_{-}\rtimes L$. Then we let $E_{\mu}$, defined as 
\[
	E_{\mu}\coloneqq \lbrace (lu_{+},l^{(q)}u_{-})\mid l\in L, u_{+}\in U_{+}, u_{-}\in U_{-}^{(q)} \rbrace\subseteq P_{+}\times P_{-}^{(q)},
\]
act on $G$ via $(e,e')\cdot g\coloneqq ege'^{-1}$. The quotient stack $ \Zip{G}^{\mu} = \quot{G/E_{\mu}}$ is the stack of $G$-zips of type $\mu$ of \cite{PWZ}.

In \cite[Cor. 6.27]{VW} they show that there exists a universal homeomorphism 
\[
	\phi\colon \nSht{1}_{G}^{\mu}\to \Zip{G}^{\mu}.
\]
In the following, we will need a slight extension of this result in the sense that the universal homeomorphism $\phi$ is actually compatible with the quotient structure:

\begin{lem}
\label{lem.univ.hom}
	There exists a universal homeomorphism $E_{1}(G,\mu)\to E_{\mu}$ that is compatible with the respective actions on $G$ inducing $\phi$.
\end{lem}
\begin{proof}
	The group $E_{1}(G,\mu)$ is given by $(U_{+}\times U_{-})\rtimes L$ \cite[Prop. 6.13]{VW}.  The action of $E_{1}(G,\mu)$ on $G$ is given by $(e,e')\cdot g \coloneqq eg(e'^{(q)})^{-1}$. 
	Note that $E_{\mu}$ can be rewritten as $(U_{+}\times U_{-}^{(q)})\rtimes L$ where the action of $L$ on $U_{-}^{(q)}$ is induced by the $q$-Frobenius. 
	The action on $G$ is given by $(e,e')\cdot g\coloneqq ege'^{-1}$. Therefore, the relative Frobenius $P_{-}\rightarrow P_{-}^{(q)}$ induces a group scheme homomorphism
	$$
		E_{1}(G,\mu)\rightarrow E_{\mu}
	$$
	that is compatible with the actions on $G$ and is a universal homeomorphism \cite[0CCB]{stacks-project}.
\end{proof}

\section{Motives on pro- and ind-Artin stacks}
In the following we want to understand the motive of $\GSht$ representing Borel-Moore homology. Recall, that each of the $\GSht^{\mu}$ is a quotient of an ind-scheme by a pro-algebraic group scheme. However, as shown in \cite{VW}, the truncated versions $\nSht{N}_{G}^{\mu}$ are in fact Artin stacks. Thus to understand motives on $\nSht{\infty}_{G}^{\mu}$ and therefore on $\Sht_{G}^{\mu}$, we need to understand motives on cofiltered systems of Artin stacks. We call these types of cofiltered system of Artin stacks \textit{pro-algebraic} (we will give a  precise definition below). 

\subsection{Recollection on the theory of motives}

For ind-Artin stacks Richarz-Scholbach defined an $\infty$-category of rational motives via right Kan extension of the classical theory on schemes \cite{RS1}.
In \cite{MotPro} the author has shown that for pro-algebraic stacks, we can define another $\infty$-category rational motives, which we will recall below, via a colimit process. For ind-Artin stacks that also admit a presentation as a pro-algebraic stack this yields a priori two different notions of the $\infty$-category of motives. But, as we will see, in special case like $\nSht{\infty}_{G}^{\mu}$ both definitions of $\DM$ become equivalent and we can benefit from both formalisms.

The benefit of the construction in \cite{MotPro} is that there is, under some assumptions, still a six functor formalism (see Theorem \ref{thm.6.ff}). In \cite{RS1} the existence of a $6$-functor formalism is only shown for ind-schemes of ind-(finite type). However, the insights of \cite{MotPro} enable us to obtain a six functor formalism for rational motives on ind-Artin stacks under the usual finiteness assumptions. In fact, we will see that the existence of a localization sequence can be shown without any finiteness assumption. This is a key insight that enables us to compute motives of stratified stacks (e.g. $\GSht$ which does not admit a covering by ind-schemes of ind-(finite type)) by the motives of the strata. \par

Let us remark here that we do need both formalisms of \cite{RS1} and \cite{MotPro}. The stack $\GSht$ is a priori just a stack for the fpqc-toplogy. However, the explicit construction as a quotient reduces to the study of motives on the ind-scheme $LG$ which is not if ind-(finite type) and motives on $L^{+}G$, which is a pro-scheme. Moreover, the stack $\nSht{\infty}_{G}$, which we will use to compute the motive of $\GSht$, is an honest pro-algebraic stack.

\medskip
In the following we recall some aspects of rational motives needed in this article, following \cite{CD}, \cite{RS1}, and \cite{MotPro}.\par
Let $p>0$ be a prime and let $S$ be a regular scheme of finite type over $\FF_{p}$.
For any scheme $X$ of finite type over $S$, denote by $\DM(X)$ the $\infty$-category of Beilinson motives over $X$ \cite{CD}. The assignment $X\mapsto \DM(X)$, $f\mapsto f^{*}$ satisfies h-descent and forms part of a six-functor formalism, due essentially to Ayoub \cite{AyoubThesis} and in the $\infty$-categorical setting to Robalo \cite{Robalo}. We will not recall details here, as this is classical and beyond the scope of this summary, but we emphasize two relevant facts about $\DM$.

Since $\DM$ satisfies \'etale descent, it extends to Artin stacks via gluing along a smooth atlas. Note, however, that the comparison between motivic cohomology and $K$-theory fails in the equivariant setting, though there remains a comparison with Edidin–Graham’s equivariant Chow groups \cite[Thm.~2.2.10]{RS1}.

Richarz–Scholbach show that this construction yields an $\infty$-category of rational motives for ind-Artin stacks locally of finite type over $S$ \cite{RS1}. In this case one must distinguish between two gluings: using $f^{*}$ or $f^{!}$. For ind-Artin stacks these differ in general. We write $\DM^{*}(X)$ for the category obtained by gluing via $f^{*}$ and $\DM^{!}(X)$ for that obtained via $f^{!}$. For technical reasons, Richarz–Scholbach work with $\DM^{!}$. Moreover, the six functors extend in the setting of ind-schemes under suitable assumptions (cf.~\emph{loc.\ cit.}), and an \'etale realization functor still exists \cite[Thm.~2.3.7]{RS1}.

Coming back to pro-systems of Artin stacks, we will summarize the necessary definitions and the main result of \cite{MotPro} for motives on pro-systems of Artin stacks.

\begin{defi}[\protect{\cite{MotPro}}]
\label{defi.p-stack}
Let $x\colon \NN^{\op}\rightarrow \PreStk_{S}$ be a diagram of prestacks over $S$ and let $X\coloneqq \lim x$ be the associated limit prestack. We call the tuple $(x,X)$ a \textit{pro-algebraic stack}, if $x(i)$ is an algebraic stack locally of finite type over $S$ for all $n\in \NN$.\par 
	If $(x,X)$ is a pro-algebraic stack, then we denote by $X_{n}\coloneqq x(n)$ and the transition maps $X_{m}\rightarrow X_{n}$ by $x_{nm}$, for $n\leq m\in\NN$.\par
	There is an obvious notion of a morphisms and properties.
\end{defi}

\begin{example}
	The assignment $\tau\colon N\mapsto \nSht{N}_{G}$ yields by design a pro-algebraic stack. In \cite{MotPro} the functor $\tau\colon N\mapsto \mathrm{Disp}_{N}$ was considered and motives on $(\tau,\Disp)$ were studied, where $\mathrm{Disp}_{N}$ denotes the stack of $N$-truncated displays of Lau \cite{LauDisp}. 
\end{example}
 
 \begin{defi}[\protect{\cite{MotPro}}]
 	Let $(x,X)$ be a pro-algebraic stack. Then we say that $(x,X)$ is
	\begin{enumerate}
		 \item \textit{classical} if $x_{nm}$ is affine for all $n\leq m\in \NN$,
		\item \textit{tame} if $x_{nm}$ is smooth for all $n\leq m\in \NN$, and
		\item \textit{strict} if $x_{nm}^{*}\colon \DM^{*}(X_{n})\rightarrow \DM^{*}(X_{m})$ is fully faithful for all $n\leq m\in \NN$.
	\end{enumerate}
 \end{defi}

Let us note that the property \textit{classical} yields representability of the limit and strictness ensures that the motivic cohomology of the limit $X$ agrees with the motivic cohomology of $X_{1}$.

\begin{lem}
\label{lem-strict}
	Let $G$ be a split reductive group and $\mu\in X_{*}(T)_{+}$ a cocharacter. The pro-algebraic stack $(\tau, \nSht{\infty}_{G}^{\mu})$ is strict tame and thus $(\tau,\nSht{\infty}_{G})$ is also strict tame.
\end{lem}
\begin{proof}
	Let us first show that $(\tau, \nSht{\infty}_{G}^{\mu})$ is strict tame. The morphism $E_{N+1}(G,\mu)\to E_{N}(G,\mu)$ is surjective with kernel given by a vector group \cite[Prop. 6.16]{VW}. 
	Therefore, by $\AA^{1}$-invariance \cite[Prop. 2.2.11]{RS1}, it is enough to show that for any $N'>N$ the map $L^{(N')}(G)\rightarrow L^{(N)}(G)$ is a torsor under a unipotent group. But this follows from \cite[Prop. A.4.9]{RS1}.
	
	Let us now show how this implies that also $(\tau,\nSht{\infty}_{G})$ is strict tame. By the split assumption on $G$, we have for all $N\in\NN$ a decomposition
	$$
		\nSht{N}_{G} = \coprod_{\mu\in X_{*}(T)_{+}} \nSht{N}_{G}^{\mu}
	$$
	\cite{VW}. This reduces us to the first case, which we already proved. 
\end{proof} 

For a pro-algebraic stack $(x,X)$, we construct a DG-category of motives on $(x,X)$ via the formula
$$
	\DM^{*}(x,X)\coloneqq \colim_{n\in \NN, x^{*}} \DM^{*}(X_{n})
$$
where we take the colimit in $\mathrm{Mod}_{\Dcal(\QQ)}(\Pr^{L})$.
This assignment can be made functorial and in \cite{MotPro} the author shows that we still obtain a $6$ functor formalism, which is motivic in a suitable sense.

\begin{thm}[\protect{\cite[Thm. 5]{MotPro}}]
\label{thm.6.ff}
	Let $(x,X)$ be a pro-algebraic stack. Then $\DM^{*}(x,X)$ is a compactly generated  closed symmetric monoidal $\infty$-category. The assignment 
	$$
	(x,X)\mapsto \DM^{*}(x,X),\quad f\mapsto f^{*}
	$$
	is part of a $6$-functor formalism $(f^{*}\dashv f_{*}, f_{!}\dashv f^{!},\otimes\dashv \Homline)$ on $!$-adjointable\footnote{See Example \ref{ex-adjointable}.} morphisms locally finite type. If one restricts to morphisms, where each square is cartesian, then this $6$-functor formalism is also motivic and orientable in the sense that it admits an orientation satisfies homotopy invariance, $T$-stability, localization, purity and Tate-twists compute Thom-motives.
\end{thm}

\begin{example}
\label{ex-adjointable}
	The condition $!$-adjointable is satisfied in all of the cases that interest us. For example if $(x,X)$ is strict tame and $(y,Y)$ is the constant diagram. In particular, for the purpose of this article the reader may forget about this condition, as it will be satisfied by the objects of interest.
\end{example}

While for ind-Artin stacks, we have the ambiguity that there are two possible definitions of $\DM$ via $f\mapsto f^{*}$ or $f\mapsto f^{!}$, this is not the case for tame pro-algebraic stacks. 

\begin{prop}[\protect{\cite[Prop. 3.13]{MotPro}}]
\label{prop-shriek}
	Let $(x,X)$ be a tame pro-algebraic stack. Then we have an equivalence 
	$$
		\DM^{*}(x,X)\simeq \colim_{n\in \NN, x^{!}} \DM^{!}(X_{n}).
	$$
\end{prop}

In a similar fashion classical pro-algebraic stacks $(x,X)$ have representable limits. This allows a continuity result, which connects the two definitions $\DM^{*}(x,X)$ and $\DM^{*}(X)$.

\begin{prop}[\protect{\cite[Prop. 2.13]{MotPro}}]
	\label{prop-classical}
	Let $(x,X)$ be a classical pro-algebraic stack. Then we have an equivalence 
	\[
		\DM^{*}(x,X) \simeq \DM^{*}(X).
	\]
\end{prop}

Finally, let us define the notion of a motive. We will only need the motive representing Borel-Moore homology.

\begin{defi}
\label{def-motive}
	Let $(x,X)$ be a strict tame pro-algebraic stack over $S$ with structure map $f\colon (x,X)\rightarrow S$, where we see $S$ as a constant diagram. Then we define 
	$$
		M^{c}_{S}(x,X)\coloneqq f_{*}f^{!}1_{S}.
	$$
	If the base is clear, then we may omit the subscript $S$.\par 
	If $X$ is representable by an algebraic stack locally of finite type, we will use the notation $M^{c}_{S}(X)$ to denote the associated motives in $\DM^{*}(X)$.
\end{defi}

Note that $M^{c}_{S}(X)$ represents Borel-Moore homology (or dually cohomology with compact support).

 \subsection{Complements on motives on ind-pro-Artin stacks}
 \label{sec-complement}
 Recall the notation of Section \ref{sec-GSht}.
 In this short subsection, we want to construct for the structure map $p\colon \GSht\to \Spec(k)$ a $!$-pullback and a $*$-adjunction. Moreover, we will need later on a localization sequence, to describe $p_{*}p^{!}1_{k}$ as a decomposition of $M^{c}$ of the strata. Note that we cannot guarantee the existence of $p_{!}$ since the loop group $LG$ is not of ind-(finite type) and so $\GSht$ is not even locally of ind-(finite type) in any suitable sense. This can be avoided once we extend the theory using methods of \cite[App. C]{Hoyois}. However, we will not make use of this extension, since we will not need it for our application.
  
\vspace{6pt} We begin with the following observations. Let $X$ be an (ordinary) Artin stack over $k$, then $\DM^{*}(X)$ is equivalent\footnote{The idea is the same as in \cite{Scholze6ff}. One checks that $(p_{\bullet}^{!})$ admits a left adjoint and the counit is given by $\colim_{[n]\in\Delta} p_{n!}p_{n}^{!}\rightarrow \id$ which by purity and descent for $\DM$ is an equivalence.} to $\DM^{!}(X)$ via $(p_{\bullet}^{!})\colon \DM^{*}(X)\to \lim_{\Delta}\DM^{!}(U^{/X}_{\bullet})$, where $p\colon U\to X$ is a smooth cover and $U^{/X}_{\bullet}$ denotes the \v{C}ech nerve. It is clear from the construction that under $\varphi_{X}\colon \DM^{*}(X)\simeq \DM^{!}(X)$ for any map $f\colon X\rightarrow Y$ locally of finite type, we have 
 \begin{equation}\label{eq.f!}
	 f^{!}\simeq \varphi^{-1}_{Y}f^{*}\varphi_{X}	
 \end{equation}

 Now let us show how to construct the $*$-adjunction for any morphism of ind-schemes. Our methods do not use that we have ind-schemes at all, so instead we will even show the existence of the $*$-adjunction for morphisms of ind-Artin stacks (again without any finiteness assumption).

Let $X=\colim_{i} X_{i}$ be an ind-Artin stack with transition maps $t_{ij}$.  Since each $t_{ij}$ is a closed immersion (\'etale locally), we have $t_{ij*}\simeq t_{ij!}$ and thus an equivalence 
\begin{equation}\label{eq.colim}
	 	\colim_{t_{ij!}}\DM^{!}(X_{i})\simeq \colim_{t_{ij*}}\DM^{*}(X_{i})
\end{equation}

Consider a morphism $f\colon X\to Y$ of ind-Artin stacks (not necessarily locally of ind-(finite type)). By functoriality and (\ref{eq.colim}) we obtain for any $f\colon X\to Y$ a $*$-pushforward $f_{*}\colon \DM^{!}(X)\to \DM^{!}(Y)$, induced by $f_{i*}$ on each level. Moreover, since $t_{ij}$ are closed immersions, the induced square
\[
 	\begin{tikzcd}
		\DM^{*}(X_{i})\arrow[r,"t^{X}_{ij*}"]\arrow[d,"f_{i*}"]& \DM^{*}(X_{j})\arrow[d,"f_{j*}"]\\
		\DM^{*}(Y_{i})\arrow[r,"t^{Y}_{ij*}"]&\DM^{*}(Y_{j*})
	\end{tikzcd}
\]
 is left adjointable, i.e. the exchange map 
\[
 	f_{i}^{*}t^{Y}_{ij*}\to t^{X}_{ij*}f_{j}^{*}
\]
 is an equivalence by proper base change. Since also each $t_{i}$ is a colimit preserving functor, we can use\footnote{Note that in \textit{loc. cit.} the colimit along $*$-pullbacks is considered, so one has to be careful with the notation.} \cite[Lem. 3.9]{MotPro} to see that $f_{*}$ admits a left adjoint 
 \[
 f^{*}\colon \colim_{t^{Y}_{ij*}} \DM^{*}(Y_{i})\to \colim_{t^{X}_{ij*}}\DM^{*}(X_{i}).
 \] 
Hence, via (\ref{eq.colim}), we obtain a left adjoint $f^{*}\colon \DM^{!}(Y)\to\DM^{!}(X)$ of $f_{*}$.

 \begin{prop}
 	Let $f\colon X\rightarrow Y$ be a morphism of ind-Artin stacks. Then there exists a functor $f_{*}\colon \DM^{!}(X)\to\DM^{!}(Y)$ that preserves compact objects with the following properties.
	\begin{enumerate}
		\item The functor $f_{*}$ admits a left adjoint $f^{*}$.
		\item If $f$ is ind-proper, then $f_{*}$ is left adjoint to $f^{!}$.
		\item If $f$ is an open immersion, then $f^{!}\simeq f^{*}$.
	\end{enumerate}
 \end{prop}
 \begin{proof}
 	See discussion above for $1$. The discussion above and (\ref{eq.f!}) yield the remaining points.
\end{proof}
% 
% \begin{lem}
% 	Let $X$ be a ind-Artin stack of ind-finite type over $S$ \hl{NEEDED?}. Further, let $i\colon Z\hookrightarrow X$ be a closed immersion of ind-Artin stacks with open complement $j\colon U\hookrightarrow X$. Then we have the following (co)fiber sequences
%	\begin{align*}
%		i_{!}i^{!}&\to \id \to j_{*}j^{*},\\
%		j_{!}j^{!}&\to \id \to i_{*}i^{*}. 
%	\end{align*}
% \end{lem}
% \begin{proof}
% 	Each of the functors preserve compact objects. So, we can check this on compacts, where we use that any compact of $M\in\DM^{*}(X)$ comes from some finite level $M_{i}\in \DM^{*}(X_{i})$. In particular, the assertion reduces to Artin stacks, where it is known by descent.
% \end{proof}
 
 \begin{rem}
Since every compact on an ind-Artin stack comes from some finite level, we can prove localization sequence for ind-Artin stacks without any finiteness assumption. Under finiteness assumptions or using \cite[App. C]{Hoyois}, we obtain also a lower $!$-functor and can prove the $!$/$*$-exchange equivalence and projection formula similarly. Moreover, analogously to the proof of \cite[Thm. 2.4.2]{RS1} we have an adjunction $\otimes\dashv \Homline$. In fact, by our methods we obtain all the six functors for any morphism of ind-Artin stacks. This was previously not known, as far as the author knows.  However as remarked in \textit{loc. cit.} in general there is no monoidal unit in $\DM^{!}(X)$ for and ind-Artin stack $X$ (in fact not even if $X$ is representable by an ind-scheme). 
 \end{rem}

 Now assume that there exists a pro-algebraic group $G=\lim_{n\in \NN} G_{n}$ over $k$ with split pro-unipotent kernel acting on an ind-scheme $X= \colim_{i}X_{i}$ (not necessarily of ind-(finite-type)) over $k$.  Then let us look at the \'etale stack quotient $p\colon \quot{X/G}\to \Spec(k)$. The quotient map $X\to \quot{X/G}$ is an effective epimorphism in the \'etale topology. Thus, by descent we can compute 
 \[
 \begin{tikzcd}
 	 \DM^{!}(X/G)\simeq \lim (\DM(X)\arrow[r,"",shift right= 0.3em] \arrow[r,"",shift left= 0.3em] &\DM^{!}(X\times G) \arrow[r,"",shift right= 0.5em]\arrow[r,""] \arrow[r,"",shift left= 0.5em]&\dots ).
 \end{tikzcd}
 \] 
 via the limit along the \v{C}ech nerve of $X\to \quot{X/G}$. To construct the adjunction $p^{*}\dashv p_{*}$ and the localization sequence it is enough to glue the functors along the \v{C}ech nerve. We have already discussed that on $\DM(X)$ there is already a $*$-adjunction. So let us handle the $\DM(X\times G^{m})$ for $m\in \NN$. We have 
\begin{equation}
\label{eq.DM!}
	\DM^{!}(X\times G^{m})\simeq \colim_{(t_{ij}\times \id_{G^{m}})_{!}}\DM^{!}(X_{i}\times G^{m})
\end{equation}
noting that $G$ is in fact a faithfully flat group scheme over $k$, so $t_{ij}\times \id_{G^{m}}$ is a closed immersion and its $!$-pushforward makes sense. Since $G$ has split pro-unipotent kernel it is in particular classical and tame by our definition. Therefore, we have 
\begin{equation}
\label{eq.DM*}
	\DM^{!}(X_{i}\times G^{m})\simeq \colim_{n\in \NN,p_{n}^{!}}\DM^{!}(X_{i}\times G^{m}_{n})\simeq \colim_{n\in \NN,p_{n}^{*}}\DM^{*}(X_{i}\times G^{m}_{n})
\end{equation}
by Proposition \ref{prop-shriek} and \ref{prop-classical}, where $p_{n}\colon G^{m}_{n+1}\to G^{m}_{n}$ denotes the canonical projection. Similar to our discussion on the $*$-adjunction on ind-Artin stacks, equivalences (\ref{eq.DM!}) and (\ref{eq.DM*}) yield an adjunction $p^{*}\dashv p_{*}$ and a localization sequence, allowing us to compute $M^{c}(\GSht)$ later on.

\begin{lem}
\label{lem.loc.seq}
	Let $X$ be an ind-scheme over $k$ and $i\colon Y\hookrightarrow X$ a closed immersion with complement $j\colon U\hookrightarrow X$. Moreover, let $G$ be a  pro-algebraic group with split pro-unipotent kernel acting on $X$. Assume further that the action of $G$ restricts to an action on $Y$ and $U$ such that $i$ and $j$ are $G$-equivariant. Let $p^{X}\colon X\rightarrow \Spec(k)$ denote the structure map and similarly for $Y$ and $U$ we denote by $p^{Y}$ and $p^{U}$ the structure maps. Then for $M\in\DM(k)$ the following sequence
	\[
		p^{Y}_{*}p^{Y!}M\to p^{X}_{*}p^{X!}M\to p^{U}_{*}p^{U!}M
	\]
	is a fiber sequence for any $\DM(k)$.
\end{lem}
\begin{proof}
	See discussion above.
\end{proof}

\section{Motives on local $G$-shtukas}
 
 Before continuing with the proof of the main theorem let us a conceptual lemma addressing the potential ambiguity between the definition of $\DM$ for prestacks via Kan extensions and $\DM$ for pro-algebraic stacks via colimits.

\begin{lem}
\label{lem-mot-kan}
	Let $S$ be a scheme.
 	We fix a pro-system $X_{\bullet}\colon \NN^{\op}\to \Sch^{\mathrm{lft}}_{S}$ of locally of finite $S$-schemes. Moreover, we assume that each $X_{m}\to X_{n}$ for $m\geq n$ is an affine bundle. 
	By definition, we set $X_{\infty}\coloneqq \lim_{n}X_{n}$.
	
	We further fix a pro-algebraic group scheme $G_{\infty} = \lim_{n} G_{n}$ with split pro-unipotent kernel such that each $G_{n}$ acts on $X_{n}$ and the transition maps $X_{m}\to X_{n}$ are equivariant with respect to $G_{m}\to G_{n}$. 
	
	The action of $G_{\bullet}$ on $X_{\bullet}$ yields a pro-algebraic stack $(x,\quot{X_{\infty}/G_{\infty}})$.
	
	\medskip
	
	The canonical map $\DM^{*}(x,\quot{X_{\infty}/G_{\infty}})\rightarrow  \DM^{!}(\quot{X_{\infty}/G_{\infty}})$ is an equivalence.
\end{lem}
\begin{proof}
	First note that $\DM^{*}(x,\quot{X_{n}/G_{n}})\simeq \DM^{!}(x,\quot{X_{n}/G_{n}})$ by Proposition \ref{prop-shriek}. The rest of the proof is completely analogous to the proof of \cite[Prop. 4.15]{MotPro}
\end{proof}

 \subsection{Application to the stack of local $G$-shtukas}
 
We will use the notation of Section \ref{sec-GSht}: Let $k=\FF_{q}$ be a finite field of characteristic $p>0$. Let $G$ be a split reductive group scheme over $k[\![z]\!]$ with split maximal torus $T$. 

In this section, we want to show that the compactly supported cohomology of $\GSht$ can be computed by the compactly supported cohomology of $\Zip{G}$, which equivalently is given by the Chow group of $\Zip{G}$ (since $\Zip{G}$ is smooth and $0$-dimensional). 

First, we show the statement for $\nSht{1}_{G}$. Then we will use the presentation of $\nSht{\infty}_{G}^{\mu}$ as a pro-algebraic stack to show that $M^{c}(\nSht{\infty}_{G}^{\mu})\simeq M^{c}(\Zip{G}^{\mu})$. Afterwards, we will compute $M^{c}(\GSht)$ using the stratification $\coprod_{\mu} \GSht^{\mu}\rightarrow \GSht$ using that \'etale locally on $k$, we have $\GSht^{\mu}\simeq \nSht{\infty}_{G}^{\mu}$.

\begin{rem}
	Let us note that $M^{c}(\Zip{G}^{\mu})$ is a Tate motive in $\DM(k)$ \cite[Ex. 5.6]{MotGZip}, i.e. contained in the full stable cocomplete subcategory of $\DM(k)$ generated by Tate twists $1_{k}(n)$. In our specific situation it is known that the subcategory of Tate motives admits a $t$-structure realizing to the perverse $t$-structure on $\ell$-adic cohomology \cite{Levine-t}. 
\end{rem}

\begin{lem}
\label{lem-zip-sht}
	For all $\mu\in X_{*}(T)_{+}$ the $*$-pullback functor 
	$$
		\DM^{*}(\nSht{1}_{G}^{\mu})\rightarrow \DM^{*}(\Zip{G}^{\mu})
	$$ 
	is an equivalence. In particular, there exists a fully faithful functor $\DM^{*}(\Zip{G}^{\mu})\rightarrow \DM^{*}(\tau,\nSht{\infty}_{G}^{\mu})$.
\end{lem}
\begin{proof}
Recall that by Lemma \ref{lem.univ.hom} there exists a universal homeomorphism $\phi\colon \nSht{1}_{G}^{\mu}\to \Zip{G}^{\mu}$ that is compatible with the \v{C}ech nerve of the canonical atlas given by $G$ on both sides.
	 Therefore it induces a map of cosimplicial diagrams
	$$
		\begin{tikzcd}
			 \dots\arrow[r,""]\arrow[d,"",shift left = 0.5em]\arrow[d,"",shift right = 0.5em]\arrow[d,""]&\dots\arrow[d,"",shift left = 0.5em]\arrow[d,"",shift right = 0.5em]\arrow[d,""]\\
			 E_{1}(G,\mu)\times G \arrow[r,""]\arrow[d,"",shift left = 0.3em]\arrow[d,"",shift right = 0.3em]& E_{\mu}\times G\arrow[d,"",shift left = 0.3em]\arrow[d,"",shift right = 0.3em]\\
			 G\arrow[r,"\id_{G}"]& G,
		\end{tikzcd}
	$$
	which on each level is a universal homeomorphism. This induces a map of simplicial diagrams
		$$
		\begin{tikzcd}
			 \dots&\dots\arrow[l,""]\\
			 \DM^{*}(E_{1}(G,\mu)\times G)\arrow[u,"",shift left = 0.5em]\arrow[u,"",shift right = 0.5em]\arrow[u,""]& \DM^{*}(E_{\mu}\times G)\arrow[l,""]\arrow[u,"",shift left = 0.5em]\arrow[u,"",shift right = 0.5em]\arrow[u,""]\\
			 \DM^{*}(G)\arrow[u,"",shift left = 0.3em]\arrow[u,"",shift right = 0.3em]& \DM^{*}(G)\arrow[u,"",shift left = 0.3em]\arrow[u,"",shift right = 0.3em]\arrow[l,""],
		\end{tikzcd}
	$$
	that is an equivalence on each level by qfh-descent \cite[Prop. 3.2.5]{Voev}. Hence, we get an equivalence on the limit 
	$$
		\phi^{*}\colon \DM^{*}(\Zip{G}^{\mu})\rightarrow\DM^{*}(\nSht{1}_{G}^{\mu})
	$$
	\cite[Thm. 2.2.16]{RS1}.
\end{proof}

\begin{lem}
\label{lem-infty-decomp}
	The motive $M^{c}(\tau,\nSht{\infty}_{G}^{\mu})$ is equivalent to $M^{c}(\Zip{G}^{\mu})$.	Moreover, we have 
	$$
		M^{c}(\tau,\nSht{\infty}_{G})\simeq \bigoplus_{\mu\in X_{*}(T)_{+}} M^{c}(\Zip{G}^{\mu}).
	$$	
\end{lem}
\begin{proof}
	By Remark \ref{rem-Sht-coprod}, we have
	$$
		M^{c}(\nSht{1}_{G})\simeq \bigoplus_{\mu\in X_{*}(T)_{+}}M^{c}(\nSht{1}_{G}^{\mu}).
	$$
	Now the result follows from Lemma \ref{lem-strict}, Lemma \ref{lem-mot-kan}, Lemma \ref{lem-zip-sht} and \cite[Lem. 4.4]{MotPro}.
\end{proof}

\begin{thm}
\label{thm.main}
	Let $\mu\in X_{*}(T)_{+}$, then 
	$$
		M^{c}(\GSht^{\leq\mu}) \simeq \bigoplus_{\mu'\leq \mu} M^{c}(\Zip{G}^{\mu'}) .
	$$
	In particular, we have 
	$$
		M^{c}(\GSht) \simeq M^{c}(\nSht{\infty}_{G}).
	$$
\end{thm}
\begin{proof}
Let us consider the stratification 
$$
	\coprod_{\mu\in X_{*}(T)_{+}}\nSht{\infty}_{G}^{\mu}\rightarrow \GSht
$$
obtained via the discussion in Section \ref{sec-GSht} and Remark \ref{rem-infty}.
	On $X_{*}(T)_{+}$ there exists a length function $l\colon X_{*}(T)_{+}\rightarrow \ZZ_{\geq 0}$ inducing the dominance order \cite[Ex. 4.2.16]{RS1}. We will prove the result by induction over $l(\mu)$. For $l(\mu) = 0$ the result holds by design. Now let $l(\mu) = n$ be arbitrary. We have a closed immersion 
	$$
		X_{n-1}\coloneqq\GSht^{\leq n-1}\subseteq X_{n}\coloneqq \GSht^{\leq n}
	$$
	with open complement $U\coloneqq \coprod_{l(\mu) = n}\nSht{\infty}_{G}^{\mu}$. By Lemma \ref{lem.loc.seq}, we have a fiber sequence
	$$
		M^{c}(X_{n-1})\rightarrow M^{c}(X_{n})\rightarrow M^{c}(U). 
	$$
	By induction it suffices to show that the fiber sequence above splits. Thus, we are reduced to show that 
	$$
		\Hom(M^{c}(\nSht{\infty}_{G}^{\mu}), M^{c}(\nSht{\infty}_{G}^{\mu'})[1]) = 0.
	$$
	for all $\mu,\mu'\in X_{*}(T)_{+}$ such that $l(\mu) = n$ and $l(\mu')<n$.  By Lemma \ref{lem-zip-sht} and Lemma \ref{lem-mot-kan}, we have
	$$
		\Hom(M^{c}(\nSht{\infty}_{G}^{\mu}), M^{c}(\nSht{\infty}_{G}^{\mu'})[1]) = \Hom(M^{c}(\Zip{G}^{\mu}), M^{c}(\Zip{G}^{\mu'})[1]).
	$$
	As $\Zip{G}^{\mu'}$ is smooth of dimension $0$, we obtain $M^{c}(\Zip{G}^{\mu'})\simeq f_{*}1_{\Zip{G}^{\mu'}}$, where $f\colon \Zip{G}^{\mu'}\rightarrow \Spec(\FF_{q})$ is the structure map. Hence, by smooth base change it suffices to show that
	$$
		A^{0}(\Zip{G}^{\mu}\times \Zip{G}^{\mu'},-1) =0.
	$$
	But this follows from the construction of higher equivariant Chow-groups.
\end{proof}

\begin{cor}
	Assume that $G$ is not necessarily split. Then we have 
	\[
		M^{c}(\GSht)\simeq  M^{c}(\nSht{1}_{G}).
	\] 
\end{cor}
\begin{proof}
	By Lemma \ref{lem-strict} it is enough to check that $M^{c}(\nSht{\infty}_{G})\to M^{c}(\GSht)$ is an equivalence. However, this can be checked \'etale locally, which follows from Lemma \ref{lem-infty-decomp} and Theorem \ref{thm.main}.
\end{proof}

\begin{rem}
	Note that if $G$ is non-split, it is still quasi-split. We can use all the constructions of \cite{VW} but we have to be careful with the stratifications by dominant cocharacters. A priori we have to work as in \textit{op.\! cit.} with conjugacy classes of cocharacters and obtain a stratification of $\GSht$ by $\nSht{\infty}_{G}^{\mu}$ only after passage to some split extension of $k$. Thus, the main obstruction existence of a stratification $\coprod \GSht^{\mu}\to \GSht$ which to our knowledge requires the existence of a split maximal torus. 
\end{rem}

%------------------------------------------------------------------

%------------------------------------------------------------------

%==================================================================
\addcontentsline{toc}{section}{References}
\bibliographystyle{plainurl}
\bibliography{MotOnSht}

\end{document}